OpenStudy (anonymous):

Am I missing something or is this really easy, seems too easy. When the input is 4, the output of f(x) = x + 21 is ___ 25..?

OpenStudy (♪chibiterasu):

The input is what goes into 'x'. f(4) = 4 + 21 So yes, 25.

OpenStudy (♪chibiterasu):

Pretty simple, yeah.

OpenStudy (anonymous):

Evaluate the function f(p) = p2 + 3p + 1 for p = -2. -1?